OpenStudy (anonymous):

A chemist collects 156 mL of N2 gas at 25ºC and 7.9 x 10^2 torr. The volume of the gas at STP would be what?

OpenStudy (anonymous):

pV=nRT --> n= pV/RT and then that n you put in: n=V/Vm --> V= n*Vm

OpenStudy (anonymous):

What is Vm?

OpenStudy (anonymous):

molar volume -> 22,4 dm3 mol-1

OpenStudy (anonymous):

Why do you subtract 1?

OpenStudy (anonymous):

mol^-1

OpenStudy (anonymous):

For some reason I'm not coming out with any answer that makes sense to the problem.

OpenStudy (anonymous):

did you transfer pressure to Pa and temperature to K

OpenStudy (anonymous):

and volume to m3

OpenStudy (anonymous):

Yes. This is my equation. V2 = (1.039 atm * 156 mL * 298.15 K)/(273.15 K * 1 atm)

OpenStudy (anonymous):

??? n= pV/RT = 7,9*10^2 torr * 133,322 Pa torr-1 * 156 *10-6 m3 / 8,314 J K-1 mol-1 * 298 K = 6,63 *10^-3 mol V= n * Vm= 6,63*10^-3 mol * 22,4 dm3 mol-1 = 0,148 dm3 = 148 cm3 or 148 mL

OpenStudy (anonymous):

Ah... I was using the wrong gas law. Thank you for correcting me.

OpenStudy (anonymous):

not a problem